An object of mass 0.5 kg at rest started to fall from a height of 180 cm from the surface of earth calculate it's momentum when it reaches the earth's surface
1
Expert's answer
2019-05-27T10:48:30-0400

Solution:

Given:


"m=0.5 kg"


"h = 180 cm =1.8 m"


Initial potential energy of the object is:

"E_p=m*g*h"

Kinetic energy at the surface:


"E_k=\frac{mv^2}{2}"

According to the law of conservation of energy (assuming no air resistance):



"E_p = E_k"

So,

"mgh=\frac{mv^2}{2}"

Solving for v:

"v=\sqrt{2gh}"

Definition of momentum:


"p=mv"

So,


"p= m*v = m\sqrt{2gh}"

Calculating:


"p= 0.5\sqrt{2*9.8*1.8}\approx 2.97 \frac{kg*m}{s}"

Answer:

"p \approx 2.97 \frac{kg*m}{s}"